bob1011:

Using the transformation T: (x, y) (x + 2, y + 1), find the distance named. Find the distance CC'

justjm:

Essentially the two points you have are (x,y) and (x+2, y+1), so you can find the distance using the distance formula \(\sqrt{(x_2 - x_1)^2 + (y_2-y_1)^2}\) \(D_{CC'}=\sqrt{(x+2-x)^2 + (y+1-y)^2}\) You can feel free to simplify, and you'll get an actual numerical value.

justjm:

@bob1011 are you there?